OpenStudy (lena772):

The following function defines a recursive sequence. f(0) = -4 f(1) = 12 f(n) = -3•f(n -1) - 2•f(n - 2); for n > 1 Which of the following sequences is defined by this recursive function? -4, 12, -28, 60, … -4, -12, -28, -60, … -4, 12, -18, 54, … -4, 12, -18, -54, …

OpenStudy (the_fizicx99):

For question one are you trying to solve using f(0) and f(1) ?

OpenStudy (schrodingers_cat):

Its is -4,12,-28,60 ... f(2) = -3(12) - 2(-4) =-28 f(3) = -3(-28) - 2(12) = 60 Hope this Helps :)

OpenStudy (lena772):

tHANK YOU @Schrodingers_Cat *FOLLOWS*

hero (hero):

The ping feature doesn't work well at the moment Lena.

OpenStudy (lena772):

ah ok, post the url here?